# UltimateKalman: Flexible Kalman Filtering and Smoothing Using Orthogonal Transformations

### v2

@@ 1 Introduction
The invention of the Kalman filter by Rudolf E. Kálmán in 1960 (Kalman, 1960) is considered one of the major inventions of the 20th century. The filter efficiently and incrementally tracks the hidden state of a linear discrete dynamic system; each state estimate uses all the observations of the system up to that point in time. The filter can also predict future states and with suitable adaptations, can handle non-linear dynamic systems and smooth entire state trajectories. The literature on Kalman filters and their applications is vast and the importance of Kalman filters is beyond doubt. We mention a few relatively recent and relatively comprehensive sources  (Brown and Hwang, 1997; Grewal and Andrews, 2015; Humpherys et al., 2012), but there are numerous other authoritative sources on Kalman filtering.
Twelve years later, Duncan and Horn discovered that mathematically, the Kalman filter computes the solution to a generalized linear least squares problem (Duncan and Horn, 1972). Algorithmically and numerically, however, the Kalman filter algorithm is far from state-of-the art algorithms for linear least squares problems, including key algorithms that were invented and published in the 1950s (Givens, 1954; Householder, 1958).
@@
In 1977 when Paige and Saunders discovered and published an elegant Kalman filter algorithm based on orthogonal transformations (Paige and Saunders, 1977). Their algorithm is a specialized QR factorization and it can easily implement filtering, prediction, and smoothing.
@@
The present paper and the software that it describes, called UltimateKalman, aim to make the algorithm widely available and to highlight its advantages over other Kalman filtering and smoothing algorithms. Indeed, the Paige and Saunders algorithm is more flexible than other Kalman algorithms in two important senses. First, unlike other Kalman filters, it does not need to know the expectation of the initial state of the system. Second, it can be easily generalized to handle quantities that are added or dropped from the state vector. These features make modeling easier, as we demonstrate with concrete examples in Sections 5.4 and 5.5. The algorithm can also easily handle problems with a varying number of observations and with missing observations, and it is equally good as a filter and smoother. Many other Kalman filter algorithms lack these two characteristics, but some do posses them, so they are not completely unique.
@@ 2.1 Discrete linear dynamic systems
where $H_{i}\in\mathbb{R}^{\ell_{i}\times n_{i}}$ and $F_{i}\in\mathbb{R}^{\ell_{i}\times n_{i-1}}$ are known full-rank matrices, $c_{i}\in\mathbb{R}^{\ell_{i-1}}$ is a known vector, called a control vector, that represents external forces acting on the system, and $\epsilon_{i}$ is an unknown noise or error vector. The control vector is often assumed to be the product of a known matrix and a known vector, but this is irrelevant for the Kalman filter. The noise or error vector $\epsilon_{i}$ admits state vectors that do not satisfy the equation $H_{i}u_{i}=F_{i}u_{i-1}+c_{i}$ exactly. The matrix $H_{i}$ is often assumed to be the identity matrix, but we do not require this (and do not require it to be square). When $H_{i}$ and $F_{i}$ are square and full rank, $u_{i}$ is a function of $u_{i-1}$ and $c_{i}$ up to the error term. Obviously, the first state $u_{0}$ that we model is not defined by an evolution recurrence.
@@
where $G_{i}\in\mathbb{R}^{m_{i}\times n_{i}}$ is a known full-rank matrix, $o_{i}\in\mathbb{R}^{m_{i}}$ is a known vector of observations (measurements), and $\delta_{i}$ represents unknown measurement errors or noise. The dimension $m_{i}$ of the observation of $u_{i}$ can vary; it can be smaller than $n_{i}$ (including zero, meaning that there are not observations of $u_{i}$), equal to $n_{i}$, or greater than $n_{i}$.
@@
If $e$ is random, has zero expectation $\operatorname{E}(e)=0$, and has a known covariance matrix $\operatorname{cov}(e)=\operatorname{E}(ee^{T})$ with $\operatorname{cov}(e)^{-1}=U^{T}U$, then the solution $\hat{u}$ of the generalized least squares problem
@@ 2.2 Kalman filters and smoothers
Some Kalman filtering and smoothing algorithms make additional assumptions, but UltimateKalman requires none. (For example, many Kalman filters assume that $G_{0}=I$, which is equivalent to assuming that the expectation of $u_{0}$ is known.)
The generalized least-squares solution of Equation (3) estimates all the state vectors $u_{1},\ldots,u_{k}$ using all the observations up to and including step $k$. We denote the vectors that make up of this solution by
@@ 3 UltimateKalman and its heritage
This section describe the UltimateKalman algorithm. The algorithm is a slight simplification of the algorithm of Paige and Saunders (Paige and Saunders, 1977) in that UltimateKalman uses block orthogonal transformations whereas the algorithm of Paige and Saunders uses Givens rotations. The algorithm is also a generalization of the algorithm of Paige and Saunders, in that we allow the user to specify $H_{i}$ and we allow the dimension of the state vector to change from step to step.
@@ 3.1 A specialized QR factorization
The diagonal blocks $R_{i,i}$ are normally square and upper triangular, but are also allowed to be rectangular with more columns than rows. The superdiagonal blocks $R_{i-1,i}$ are are not identically zero. The same series of transformations is applied to the weighted right-hand side vector
@@ 3.3 The Paige-Saunders factorization algorithm
If this block is flat (or more generally, if its rank is smaller than $n_{k}$, which implies that it can be orthogonally reduced to a flat block), the algorithm leaves the bottom $2$-by-$2$ block as is, denoting
$\left[\begin{array}[]{cc}R_{k-1,k-1}&R_{k-1,k}\end{array}\right]=\left[\begin{array}[]{cc}\tilde{R}_{k-1,k-1}\\ -V_{k}F_{k}&V_{k}H_{k}\end{array}\right]\;.$
Otherwise, the algorithm computes a QR factorization of the block (9), uses the resulting $R$ factor as $R_{k-1,k-1}$, and applies the orthonormal transformation to the last block column and to the right-hand side $y$. This transforms the $R$ factor into
@@
We now have $R^{(k)}$ and the $y^{(k)}$. Note that we have not used in this step block rows $21,\ldots,k-$ 2 of $R^{(k-1)}$ and $y^{(k-1)}$.
@@ 3.4 Forgetting and rolling back
Given $R^{(k)}$ and the $y^{(k)}$, the estimates $\hat{u}_{i|k}$ are computed by back substitution, from the bottom up.
@@ 4 Implementation
The MATLAB implementation does not rely on any MATLAB toolbox, only on functionality that is part of the core product. The implementation also works under GNU Octave. The C implementation relies on basic matrix and vector operations from the BLAS (Dongarra et al., 1990b; Dongarra et al., 1990a) and on the QR and Cholesky factorizations from LAPACK (Anderson et al., 1999). The Java implementation uses the Apache Commons Math library for both basic matrix-vector operations and for the QR and Cholesky factorizations.
@@ 4.1 Programming interface
The the MATLAB and Java implementations use overloading (using the same method name more than once, with different argument lists) to express variants of these operations, while the C implementation uses NULL to express missing or default values. The estimate method returns two values in the MATLAB implementation, the estimated vector and its covariance matrix, but only the estimated vector in the Java and C implementations. The covariance matrix is returned in these implementations by a separate method, covariance. These design decisions reflect good programming practices in each programming language.
@@ 5.1 Basic tests
Figure 1: Kalman filtering, prediction, and smoothing of the trajectory of a rotating point in the plane. On the left $G=I$ and observations are indicated by red dots. The black curve represents a simulated system; the blue curve represents filtered estimates, the magenta whole-trajectory smoothed estimates, and the cyan predictions from one observation. On the right $G=\begin{bmatrix}1&0\end{bmatrix}$; we cannot predict the estate from one such observation, and we cannot produced a filtered estimate of the first state.
@@ 5.2 Variance variations
Figure 2: Kalman filtering and smoothing of a scalar trajectory. The black lines show the actual trajectory, the red dots show the observations, the blue lines show the filtered estimates (left column), the magenta lines show smoothed estimates (right column), and the gray areas represents values that are ±$\pm 3\hat{\sigma}_{i}$ from the estimate $\hat{u}_{i}$, where $\hat{\sigma}_{i}$ is the estimated standard deviation of $\hat{u}_{i}$. In the top row, the standard deviation of $\delta_{i}$ is always $10$; in the bottom row, it drops down to $0.25$ at iteration $50$ and then goes back up to $10$.
@@ 5.3 Adding and removing parameters
which represent a normal evolution of the first component of the state while not using any historical information about the second, new component. The evolution matrix $F_{2}$ remains a $1$-by-$1$ identity in this step, but $G_{2}$ is now a $2$-by-$2$ identity; $o_{2}$ becomes $2$-dimensional as well. The actual value of the second parameter is $2$ and the observations reflect that. In the next iteration, $F_{3}$ grows to $2$-by-$2$. After two iterations with a 2-dimensional state vector, we drop the first (original) component of the state vector by calling evolve with a first argument $1$ and by providing an explicit matrix
@@ 5.4 A Projectile problem
Figure 3: Kalman filtering and smoothing of the trajectory of a projectile affected by gravity and drag. Again, the black line show the actual trajectory, the red dots show the observations, the blue line shows the filtered estimates of the trajectory, and the magenta line shows the smoothed estimates. Smoothing was performed on the entire trajectory, to estimate the point of departure and the point where the projectile hits the ground. The graph on the left shows the entire trajectory and the graph on the right only the part in which observations are available.
@@ 5.5 Clock offsets in a distributed system
Figure 4: Kalman tracking of the relative offsets of three clocks from times of arrival of radio packets. The graphs show the relative offsets relative to the offset of the first (blue) clock. The solid lines are the simulated offsets and the dashed lines are the filtered estimates.
The clock_offsets example highlights the utility of the $H_{i}$ matrices. The aim is to estimate the relative offsets of a set of clocks in a distributed system, such as a wireless sensor network. Each clock is associated with a receiver and we assume that at time $\tau$, clock $j$ shows $t_{j}=\tau+f_{ij}$ where $f_{ij}$ is the offset of clock $j$ from real time at the time in which it displays the value $t_{j}$. The receivers receive radio packets from a beacon transmitter. The locations of the transmitter and receivers are known, so the line-of-sight propagation delays $d_{j}$ between the transmitter and receiver $j$ are also known. The receivers estimate the time of arrival of the packets using their local imperfect clocks. The observation equation for the time of arrival of packet $i$ at receiver $j$ is
@@ 5.6 Performance testing
Figure 5 shows the result of testing the performance of the three implementations using their perftest method. The experiments were carried out on a laptop with an Intel i7 processor. The test uses random square unitary matrices for $F_{i}$ and $G_{i}$ are fixed, identities for $C_{i}$ and $K_{i}$, $c_{i}=0$, and a random Gaussian vector for $o_{i}$. The matrices $H_{i}$ are identities created by the algorithm itself; they are not passed as arguments. The use of unitary matrices avoids overflows and underflows.
The graphs in Figure 5 show the average running times per step, averaged over groups of 1000 steps.
@@
Some of the changes in the running times that are visible in some of the plots are likely due to the computer, a laptop, slowing down to avoid overheating.
Smoothing takes more time per step, due to the back substitution phase, and perhaps more importantly, a lot more memory, because the code retains the blocks of $R$ and the transformed $b$. Still, the library can smooth very long sequences. On the same computer we were able to smooth a 5 million steps with $n_{i}=m_{i}=6$ in about 96 s and $100,000$ steps with $n_{i}=m_{i}=48$ in 84 s. Both runs required less than 16GB of memory.
@@ 6 Discussion
The Paige-Saunders algorithm does have two limitations that UltimateKalman inherits. First, it assumes that the covariance matrices $C_{i}$ and $K_{i}$ are all nonsingular. Traditional Kalman filters can cope with singular covariance matrices. However, this is not a significant limitation, since physical observations are always associated with some level of uncertainty, and physical evolution equations usually do too. Continuous evolution equations can be exact, with no uncertainty (e.g., equations of motion), but discrete evolution equations have at least discretization (truncation) error terms. Even continuous evolution equations sometimes suffer from uncertainty due to incomplete physical modeling (e.g, a missing or simplified drag term). Second, the algorithm is almost completely sequential, except for parallelism within the matrix operations that are carried out in each step. In real-time filtering applications, this limitation is meaningless, since the application is inherently sequential. However, in post-processing applications the ability to use many cores to reduce running times could be beneficial; the Paige-Saunders algorithm and UltimateKalman lack this ability.
@@
We hope that the numerical robustness afforded by the use of orthogonal transformations, the flexibility, the convenient programming interfaces, and the availability of the algorithm in multiple languages will make UltimateKalman the standard linear Kalman filter and smoother in new software. We also hope that authors and maintainers of statistical software packages of wider scope (Peng and Aston, 2011; Villegas and Pedregal, 2018), which often include nonlinear Kalman filters and other state-space models, will also incorporate UltimateKalman into their packages, ideally exposing some of the unique features through their own programming and user interfaces. Finally, we hope that authors of code generators that automatically generate Kalman filter code optimized for specific cases (Whittle and Schumann, 2004) will also incorporate the algorithm into their generators; The code clarity that we strived to achieve should simplify such efforts, as well as efforts to implement the algorithm in additional programming languages that are widely used for numerical computing like Python and Julia.

### v3

@@ 1 Introduction
The invention of the Kalman filter by Rudolf E. Kálmán in 1960 (Kalman, 1960) is considered one of the major inventions of the 20th century. The filter efficiently and incrementally tracks the hidden state of a linear discrete dynamic system; each state estimate uses all the observations of the system up to that point in time. The filter can also predict future states and with suitable adaptations, can handle nonlinear dynamic systems and smooth entire state trajectories. The literature on Kalman filters and their applications is vast and the importance of Kalman filters is beyond doubt. We mention a few relatively recent and relatively comprehensive sources  (Brown and Hwang, 1997; Grewal and Andrews, 2015; Humpherys et al., 2012), but there are numerous other authoritative sources on Kalman filtering.
Twelve years later, Duncan and Horn discovered that mathematically, the Kalman filter computes the solution to a generalized linear least squares problem (Duncan and Horn, 1972). Algorithmically and numerically, however, the Kalman filter algorithm is far from state-of-the art algorithms for linear least squares problems. In particular, it does not use orthogonal factorizations, which were invented and published in the 1950s (Givens, 1954; Householder, 1958).
@@
In 1977, Paige and Saunders discovered and published an elegant Kalman filter algorithm based on orthogonal transformations (Paige and Saunders, 1977). Their algorithm is a specialized QR factorization and it can easily implement filtering, prediction, and smoothing.
@@
The present paper and the software that it describes, called UltimateKalman, aim to make the algorithm widely available and to highlight its advantages over other Kalman filtering and smoothing algorithms. Indeed, the Paige and Saunders algorithm is more flexible than other Kalman algorithms in two important senses. First, unlike other Kalman filters, it does not need to know the expectation of the initial state of the system. Second, it can be easily generalized to handle quantities that are added or dropped from the state vector. These features make modeling easier, as we demonstrate with concrete examples in Sections 5.4 and 5.5. The algorithm can also easily handle problems with a varying number of observations and with missing observations, and it can both filter and smooth. Many other Kalman filter algorithms lack these two characteristics, but some do posses them, so they are not completely unique.
@@ 2.1 Discrete linear dynamic systems
where $H_{i}\in\mathbb{R}^{\ell_{i}\times n_{i}}$ and $F_{i}\in\mathbb{R}^{\ell_{i}\times n_{i-1}}$ are known full-rank matrices, $c_{i}\in\mathbb{R}^{\ell_{i}}$ is a known vector, called a control vector, that represents external forces acting on the system, and $\epsilon_{i}$ is an unknown noise or error vector. The control vector is often assumed to be the product of a known matrix and a known vector, but this is irrelevant for the Kalman filter. The noise or error vector $\epsilon_{i}$ admits state vectors that do not satisfy the equation $H_{i}u_{i}=F_{i}u_{i-1}+c_{i}$ exactly. The matrix $H_{i}$ is often assumed to be the identity matrix, but we do not require this (and do not require it to be square). When $H_{i}$ and $F_{i}$ are square and full-rank, $u_{i}$ is a function of $u_{i-1}$ and $c_{i}$ up to the error term. Obviously, the first state $u_{0}$ that we model is not defined by an evolution recurrence.
@@
where $G_{i}\in\mathbb{R}^{m_{i}\times n_{i}}$ is a known full-rank matrix, $o_{i}\in\mathbb{R}^{m_{i}}$ is a known vector of observations (measurements), and $\delta_{i}$ represents unknown measurement errors or noise. The dimension $m_{i}$ of the observation of $u_{i}$ can vary; it can be smaller than $n_{i}$ (including zero, meaning that there are no observations of $u_{i}$), equal to $n_{i}$, or greater than $n_{i}$.
@@
If $e$ is random, has zero expectation $\operatorname{E}(e)=0$, and has a known nonsingular covariance matrix $\operatorname{cov}(e)=\operatorname{E}(ee^{T})$ with $\operatorname{cov}(e)^{-1}=U^{T}U$, where $U$ is square then the solution $\hat{u}$ of the generalized least squares problem
@@ 2.2 Kalman filters and smoothers
Most Kalman filtering and smoothing algorithms make additional assumptions. For example, many Kalman filters assume that $G_{0}=I$, which is equivalent to assuming that the expectation of $u_{0}$ is known., UltimateKalman also makes an additional assumption: that all the $C_{i}$s and all the $K_{i}$s are nonsingular.
The generalized least-squares solution of Equation (3) estimates all the state vectors $u_{1},\ldots,u_{k}$ using all the observations up to and including step $k$. We denote the vectors that make up this solution by
@@ 3 UltimateKalman and its heritage
This section describes the UltimateKalman algorithm. The algorithm is a slight simplification of the algorithm of Paige and Saunders (Paige and Saunders, 1977) in that UltimateKalman uses block orthogonal transformations whereas the algorithm of Paige and Saunders uses Givens rotations. The algorithm is also a generalization of the algorithm of Paige and Saunders in that we allow the user to specify $H_{i}$ and we allow the dimension of the state vector to change from step to step.
@@ 3.1 A specialized QR factorization
The diagonal blocks $R_{i,i}$ are normally square and upper triangular, but are also allowed to be rectangular with more columns than rows. The superdiagonal blocks $R_{i-1,i}$ are not identically zero. The same series of transformations is applied to the weighted right-hand side vector
@@ 3.3 The Paige-Saunders factorization algorithm
If this block is flat (or more generally, if its rank is smaller than $n_{k}$, which implies that it can be orthogonally reduced to a flat block), the algorithm leaves the bottom $2$-by-$2$ block as is, but treats it as a $1$-by-$2$ block matrix $\begin{bmatrix}R_{k-1,k-1}&R_{k-1,k-1}\end{bmatrix}$, with the same column partitioning. This allows the algorithm to handle all shapes of (9) in a uniform way. That is, we rename the blocks as
$\left[\begin{array}[]{cc}R_{k-1,k-1}&R_{k-1,k-1}\end{array}\right]=\left[\begin{array}[]{cc}\tilde{R}_{k-1,k-1}\\ -V_{k}F_{k}&V_{k}H_{k}\end{array}\right]\;.$
Otherwise, the algorithm computes a QR factorization of (9), uses the resulting $R$ factor as $R_{k-1,k-1}$, and applies the orthonormal transformation to the last block column and to the right-hand side $y$. This transforms the $R$ factor into
@@
We now have $R^{(k)}$ and $y^{(k)}$. Note that we have not used in this step block rows $21,\ldots,k-$ 2 of $R^{(k-1)}$ and $y^{(k-1)}$.
@@ 3.4 Forgetting and rolling back
Given $R^{(k)}$ and $y^{(k)}$, the estimates $\hat{u}_{i|k}$ are computed by back substitution, from the bottom up.
@@ 4 Implementation
The MATLAB implementation does not rely on any MATLAB toolbox, only on functionality that is part of the core product. The implementation also works under GNU Octave. The C implementation relies on basic matrix and vector operations from the BLAS (Dongarra et al., 1990b, a) and on the QR and Cholesky factorizations from LAPACK (Anderson et al., 1999). The Java implementation uses the Apache Commons Math library for both basic matrix-vector operations and for the QR and Cholesky factorizations.
@@ 4.1 Programming interface
The MATLAB and Java implementations use overloading (using the same method name more than once, with different argument lists) to express variants of these operations, while the C implementation uses NULL to express missing or default values. The estimate method returns two values in the MATLAB implementation, the estimated vector and its covariance matrix, but only the estimated vector in the Java and C implementations. The covariance matrix is returned in these implementations by a separate method, covariance. These design decisions reflect good programming practices in each programming language.
@@ 5.1 Basic tests
Figure 1: Kalman filtering, prediction, and smoothing of the trajectory of a rotating point in the plane. On the left $G=I$ and observations are indicated by red dots. The yellow curve represents a simulated system; the blue curve represents filtered estimates, the cyan whole-trajectory smoothed estimates, and the green predictions from one observation. On the right $G=\begin{bmatrix}1&0\end{bmatrix}$; we cannot predict the state from one such observation, and we cannot produce a filtered estimate of the first state.
@@ 5.2 Variance variations
Figure 2: Kalman filtering and smoothing of a scalar trajectory. The yellow lines show the actual trajectory, the red dots show the observations, the blue lines show the filtered estimates (left column), the cyan lines show smoothed estimates (right column), and the gray areas represent values that are ±$\pm 3\hat{\sigma}_{i}$ from the estimate $\hat{u}_{i}$, where $\hat{\sigma}_{i}$ is the estimated standard deviation of $\hat{u}_{i}$. In the top row, the standard deviation of $\delta_{i}$ is always $10$; in the bottom row, it drops down to $0.25$ at iteration $50$ and then goes back up to $10$.
@@ 5.3 Adding and removing parameters
which represents a normal evolution of the first component of the state while not using any historical information about the second, new component. The evolution matrix $F_{2}$ remains a $1$-by-$1$ identity in this step, but $G_{2}$ is now a $2$-by-$2$ identity; $o_{2}$ becomes $2$-dimensional as well. The actual value of the second parameter is $2$ and the observations reflect that. In the next iteration, $F_{3}$ grows to $2$-by-$2$. After two iterations with a 2-dimensional state vector, we drop the first (original) component of the state vector by calling evolve with a first argument $1$ and by providing an explicit matrix
@@ 5.4 A Projectile problem
Figure 3: Kalman filtering and smoothing of the trajectory of a projectile affected by gravity and drag. Again, the yellow line show the actual trajectory, the red dots show the observations, the blue line shows the filtered estimates of the trajectory, and the cyan line shows the smoothed estimates. Smoothing was performed on the entire trajectory, to estimate the point of departure and the point where the projectile hits the ground. The graph on the left shows the entire trajectory and the graph on the right only the part in which observations are available.
@@ 5.5 Clock offsets in a distributed system
Figure 4: Kalman tracking of the relative offsets of three clocks from times of arrival of radio packets. The graphs show the relative offsets relative to the offset of the first clock. The yellow lines are the simulated offsets and the blue lines are the filtered estimates.
The clock_offsets example highlights the utility of the $H_{i}$ matrices. The aim is to estimate the relative offsets of a set of clocks in a distributed system. Each clock is associated with a receiver and we assume that at time $\tau$, clock $j$ shows $t_{j}=\tau+f_{ij}$ where $f_{ij}$ is the offset of clock $j$ from real time at the time in which it displays the value $t_{j}$. The receivers receive radio packets from a beacon transmitter. The locations of the transmitter and receivers are known, so the line-of-sight propagation delays $d_{j}$ between the transmitter and receiver $j$ are also known. The receivers estimate the time of arrival of the packets using their local imperfect clocks. The observation equation for the time of arrival of packet $i$ at receiver $j$ is
@@ 5.6 Performance testing
Figure 5 shows the result of testing the performance of the three implementations using their perftest method. The experiments were carried out on a laptop with an Intel i7 processor. The test uses two random square unitary matrices $F$ and $G$, sets $F_{i}=F$ and $G_{i}=G$, uses identities for $C_{i}$ and $K_{i}$, $c_{i}=0$, and a random Gaussian vector for $o_{i}$. The matrices $H_{i}$ are identities created by the algorithm itself; they are not passed as arguments. The use of unitary matrices avoids overflows and underflows.
The graphs in Figure 5 show the average runtimes per step, averaged over groups of 1000 steps.
@@
Some of the changes in the runtimes that are visible in some of the plots are likely due to the computer, a laptop, slowing down to avoid overheating.
Smoothing takes more time per step, due to the back substitution phase, and perhaps more importantly, a lot more memory, because the code retains the blocks of $R$ and the transformed $b$. Still, the library can smooth very long sequences. On the same computer we were able to smooth 5 million steps with $n_{i}=m_{i}=6$ in about 96 s and $100,000$ steps with $n_{i}=m_{i}=48$ in 84 s. Both runs required less than 16GB of memory.
@@ 6 Discussion
The Paige-Saunders algorithm does have two limitations that UltimateKalman inherits. First, it assumes that the covariance matrices $C_{i}$ and $K_{i}$ are all nonsingular. Traditional Kalman filters can cope with singular covariance matrices. However, this is not a significant limitation, since physical observations are always associated with some level of uncertainty, and physical evolution equations usually are too. Continuous evolution equations can be exact, with no uncertainty (e.g., equations of motion), but discrete evolution equations have at least discretization (truncation) error terms. Even continuous evolution equations sometimes suffer from uncertainty due to incomplete physical modeling (e.g., a missing or simplified drag term). Second, the algorithm is almost completely sequential, except for parallelism within the matrix operations that are carried out in each step. In real-time filtering applications, this limitation is meaningless, as the application is inherently sequential. However, in post-processing applications the ability to use many cores to reduce runtimes could be beneficial; the Paige-Saunders algorithm and UltimateKalman lack this ability.
@@
We hope that the numerical robustness afforded by the use of orthogonal transformations, the flexibility, the convenient programming interfaces, and the availability of the algorithm in multiple languages will make UltimateKalman the standard linear Kalman filter and smoother in new software. We also hope that authors and maintainers of statistical software packages of wider scope (Peng and Aston, 2011; Villegas and Pedregal, 2018), which often include nonlinear Kalman filters and other state-space models, will incorporate UltimateKalman into their packages, ideally exposing some of the unique features through their own programming and user interfaces. Finally, we hope that authors of code generators that automatically generate Kalman filter code optimized for specific cases (Whittle and Schumann, 2004) will also incorporate the algorithm into their generators. The code clarity that we strived to achieve should simplify such efforts, as well as efforts to implement the algorithm in additional programming languages that are widely used for numerical computing like Python and Julia.